OpenStudy (anonymous):

how to I find the inverse of y=5e^2x+1

OpenStudy (mimi_x3):

\[y=5e^{2x}+1\] to find the inverse; replace y with x \[ x = 5e^{2y}+1\] \[\ x - 1 = 5e^{2y} => 5e^{2y} = x-1 => e^{2y} = \frac{x-1}{5}\]\]

OpenStudy (dan815):